# K3 Surfaces and Orthogonal Modular Forms

### v1

@@ 1. Introduction and statement of results
Let $\mathcal{X}$ be a smooth complex projective K3 surface. Denote by $\Lambda_{K3}$ the rank-22 lattice associated with the intersection form on $H^{2}(\mathcal{X},\mathbb{Z})$. The isomorphism class of $\Lambda_{K3}$ is independent of $\mathcal{X}$. Denote by $\mathrm{NS}(\mathcal{X})$ the Néron-Severi lattice, the rank of which may vary between $1$ and $20$.
@@
(1.1) $\mathscr{D}(L)\ =\ \Big{\{}[z]\in\mathbb{P}\big{(}L\otimes\mathbb{C}\big{)}\colon\ \langle z,z\rangle=0,\ \langle z,\overline{z}\rangle>0\Big{\}}^{+}\ .$
@@
The classifying period space $\mathrm{O}^{+}(L)\backslash\mathscr{D}(L)$ may then be seen, from this point of view, as a quotient of a bounded symmetric domain by the action of a (modular) discrete group. Examples for $S$-polarizations of high rank lead to specializations of $\mathrm{O}^{+}(L)\backslash\mathscr{D}(L)$ to classical quotients in algebraic geometry: the Hilbert modular surface in rank 18 and the Siegel modular threefold of genus two in rank 17. These bounded symmetric domains were discussed in [MR1834454]*Sec. X.6.
The present paper explores the connection between algebraic descriptions for $S$-polarized K3 surfaces and orthogonal modular forms associated with the domain $\mathscr{D}(L)$ and group $\mathrm{O}(L)^{+}$. For reasons to become apparent during the exposition, we shall restrict our discussion to a specific class of lattice polarizations: these are lattices $S$ of rank 13 and higher, for which the general [^2] A $S$-polarized K3 surface $(\mathcal{X},i)$ is called general if $i(S)=\mathrm{NS}(\mathcal{X})$. $S$-polarized K3 surface $\mathcal{X}$ satisfies ${\rm Aut}(\mathcal{X})=(\mathbb{Z}/2\mathbb{Z})^{2}$. Under the above conditions, and via standard results by Nikulin [MR633160, MR556762, MR752938], Vinberg [MR2429266], and Kondo [MR1029967], one obtains that the possible polarizing lattices $S$ reduce to the following list:
@@
For the above instances, there is a good understanding of the rings of modular forms, thanks to works by Woitalla [MR3883325] and Wang and Williams [MR4130466]. Namely, consider the graded ring of modular forms on $\mathscr{D}(L)$ with respect to the group $\mathrm{O}(L)^{+}$:
(1.3) $M(\mathscr{D}(L),\mathrm{O}^{+}(L))=\bigoplus_{m=0}^{\infty}M(\mathscr{D}(L),\mathrm{O}^{+}(L))_{m},$
@@
For lattices $L$ as in the left-side column of the table below , the modular form ring $M(\mathscr{D}(L),\mathrm{O}^{+}(L))$ is freely generated by a set of modular forms with weights tabulated in the right-side column:
@@
Consider lattices $S$, as in Equation (1.2), with $\rho_{S}\geq 14$. The coefficients $\{a_{m},b_{n}\}$ or $\{c_{m},b_{n}\}$ of Equation (1.5) are modular forms with respect to the appropriate group $\mathrm{O}^{+}(L)$, and are generators for the ring $M(\mathscr{D}(L),\mathrm{O}^{+}(L))$ in the sense of Theorem 1.1.
@@
For $S=H\oplus E_{7}(-1)\oplus A_{1}(-1)^{\oplus 4}$, the coefficients $\{a_{m},b_{n}\}$ in Equation (1.5) are meromorphic modular forms on $\mathscr{D}(L)$, of weight 2, 4, 4, 6, 6, 8, 10, 12, with respect to a subgroup $\Gamma\leq\mathrm{O}^{+}(L)$ of finite co-volume. Moreover, $\{a_{m},b_{n}\}$ form a set of generators for the ring $M^{!}(\mathscr{D}(L),\Gamma)$ of meromorphic modular forms on $\mathscr{D}(L)$ which are holomorphic away from the Heegner divisor $C^{\prime}$.
@@ 4. Certain algebras of modular forms
be the basic weak Jacobi forms, which generate the $\mathbb{C}[E_{4},E_{6}]$-module of weak Jacobi forms, as defined in Chapter 9 of [EZ85].
@@
generate the $\mathbb{C}[E_{4},E_{6}]$-module of weak Jacobi forms of lattice index $A_{1}^{\oplus n}$.
@@
For $1\leq n\leq 4$, the algebra of modular forms
$M_{*}(L_{n}):=M_{*}(\mathrm{O}^{+}(L_{n}))$
is generated by the Eisenstein series $E_{4}$ and $E_{6}$ and by the theta lifts
@@
The algebra of meromorphic modular forms
$M_{*}^{!}(L_{5})=M_{*}^{!}(\mathrm{O}^{+}(L_{5}))$
with poles on $r_{2}^{\perp}$ and its orbit under $\mathrm{O}^{+}(L_{5})$ is generated by the Eisenstein series $E_{4}$ and $E_{6}$ and by the meromorphic theta lifts
@@ 5. Computation of van Geemen–Sarti equations
We now determine the coefficients of five of the van Geemen–Sarti equations (1.4) in terms of the generators of $M_{*}(L_{n})$ described in the previous section. The argument is an induction on $n=8-k$, beginning with $n=0$ (such that $K_{n}=\{0\}$ and $L=H\oplus H$); the general K3 surface in the associated family has Picard rank 18.
@@
$\mathrm{res}\colon\ M_{*}(L_{1})\longrightarrow M_{*}(L_{0})$
which through the interpretations $M_{*}(L_{1})=M_{2*}(\mathrm{Sp}_{4}(\mathbb{Z}))$ and $M_{*}(L_{0})=\mathrm{Sym}^{2}(M_{*}(\mathrm{SL}_{2}(\mathbb{Z}))$ is the map
@@
$\mathrm{res}\colon\ M_{*}(L_{2})\longrightarrow M_{*}(L_{1}).$

### v2

@@ 1. Introduction and statement of results
1.1. Introduction
Let $\mathcal{X}$ be a smooth complex projective K3 surface. The intersection form gives the second cohomology group $H^{2}(\mathcal{X},\mathbb{Z})$ the structure of an even integral lattice, i.e., a finite-rank free abelian group together with an integer-valued nondegenerate quadratic form. This lattice is the unique (up to isometry) unimodular even lattice of signature $(3,19)$, which we denote $\Lambda_{K_{3}}\cong H^{\oplus 3}\oplus E_{8}(-1)^{\oplus 2}$; in particular it is independent of $\mathcal{X}$. Denote by $\mathrm{NS}(\mathcal{X})$ the Néron–Severi lattice, the rank of which is between $1$ and $20$.
@@ 1.1. Introduction
(1.1) $\mathscr{D}(L)\ =\ \Big{\{}[z]\in\mathbb{P}\big{(}L\otimes\mathbb{C}\big{)}\colon\ \langle z,z\rangle=0,\ \langle z,\overline{z}\rangle>0\Big{\}}^{+}$
@@
The classifying period space $\mathrm{O}^{+}(L)\backslash\mathscr{D}(L)$ may then be seen, from this point of view, as a quotient of a bounded symmetric domain by the action of a (modular) discrete group.
Cases associated with specific examples for $S$-polarizations of high rank lead to specializations of the modular quotient $\mathrm{O}^{+}(L)\backslash\mathscr{D}(L)$ to some classical quotient spaces in algebraic geometry. Such cases were previously studied by the first two authors [MR2369941, MR2935386] [MR1834454]*Sec. X.6 , as well as others, and largely motivate the present work. For instance, if $S=H\oplus E_{8}(-1)\oplus E_{8}(-1)$, the modular quotient is 2-dimensional and may be identified with the Hilbert modular surface
$\left[\left(\mathrm{SL}_{2}(\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z})\right)\rtimes\mathbb{Z}/2\mathbb{Z}\right]\backslash\left(\mathbb{H}\times\mathbb{H}\right),$
where $\mathbb{H}$ is the complex upper half-plane. Similarly interesting, if $S$ is the rank-seventeen lattice $H\oplus E_{8}(-1)\oplus E_{7}(-1)$, the modular quotient in question is isomorphic to the classical Siegel modular three-fold of genus two, i.e.,
$\mathrm{Sp}_{4}(\mathbb{Z})\backslash\mathbb{H}_{2},$
where $\mathbb{H}_{2}$ is the 3-dimensional complex Siegel modular half-space.
The present paper explores the connection between algebraic descriptions for $S$-polarized K3 surfaces and orthogonal modular forms associated with the domain $\mathscr{D}(L)$ and group $\mathrm{O}(L)^{+}$. This is motivated, as mentioned earlier, by the interesting phenomena seen in the context of high-rank specific example. If $S$ is, for instance, given by $H\oplus E_{8}(-1)\oplus E_{8}(-1)$, then, as discussed in [MR2369941], all $S$-polarized K3 surfaces can be given an explicit algebraic construction via Inose forms – a special class of quartic equations in $\mathbb{P}^{3}$. Moreover, the coefficients of the Inose form can be written explicitly in terms of the $\mathrm{SL}_{2}(\mathbb{Z})$ Eisenstein series $E_{4}$ and $E_{6}$, as well as the weight-twelve discriminant $\Delta$. A similar situation arises in the case $S=H\oplus E_{8}(-1)\oplus E_{7}(-1)$. In this context (see [MR2935386]), the $S$-polarized K3 surfaces can again be described via a generalized version of Inose forms, and the Inose coefficients recover the $\mathrm{Sp}_{4}(\mathbb{Z})$ Eisenstein series $E_{4}$ and $E_{6}$, as well as the Igusa cusp forms $\chi_{10}$ and $\chi_{12}$. These four Siegel modular forms are known to generate the graded algebra of even-weight Siegel modular forms of genus two.
For the present work, we shall restrict our discussion to a specific class of lattice polarizations: those for which the general [^2] A $S$-polarized K3 surface $(\mathcal{X},i)$ is called general if $i(S)=\mathrm{NS}(\mathcal{X})$. $S$-polarized K3 surface $\mathcal{X}$ satisfies ${\rm Aut}(\mathcal{X})=(\mathbb{Z}/2\mathbb{Z})^{2}$. As we shall explain in Section 3, these polarizing lattices carry a canonical Jacobian elliptic fibration and such a fibration proves to be the essential tool for generalizing the modular form connections noticed above for specific $S$ instances of ranks $18$ and $17$. The Jacobian elliptic fibration in question can be described algebraically via a Weierstrass-type form and the polynomial coefficients appearing in this form (except for possible poles at the Noether–Lefschetz locus), turn out to be modular forms.
For reasons to be explained shortly, we consider only lattices $S$ of rank 13 and higher, Under the above conditions, and via standard results by Nikulin [MR633160, MR556762, MR752938], Vinberg [MR2429266], and Kondo [MR1029967], one obtains that the possible polarizing lattices $S$ reduce to the following list:
@@
In the above cases, there is a good understanding of the rings of modular forms, thanks to works by Woitalla [MR3883325] and Wang and Williams [MR4130466]. Namely, consider the graded ring of modular forms on $\mathscr{D}(L)$ with respect to the group $\mathrm{O}^{+}(L)$:
(1.3) $M_{*}(\mathrm{O}^{+}(L))=\bigoplus_{m=0}^{\infty}M_{m}(\mathrm{O}^{+}(L))\ ,$
@@
For lattices $L$ as in the left-side column of the table below , the modular form ring $M_{*}(\mathrm{O}^{+}(L))$ is freely generated by a set of modular forms with weights tabulated in the right-side column:
@@ 1.2. Statement of Results
Consider lattices $S$, as in Equation (1.2), with $\rho_{S}\geq 14$. The coefficients $\{a_{m},b_{n}\}$ or $\{c_{m},b_{n}\}$ of Equation (1.5) are modular forms with respect to the appropriate group $\mathrm{O}^{+}(L)$, and are generators for the ring $M_{*}(\mathrm{O}^{+}(L))$ in the sense of Theorem 1.1.
@@
For $S=H\oplus E_{7}(-1)\oplus A_{1}(-1)^{\oplus 4}$, the coefficients $\{a_{m},b_{n}\}$ in Equation (1.5) are meromorphic modular forms on $\mathscr{D}(L)$, of weight 2, 4, 4, 6, 6, 8, 10, 12, with respect to a subgroup $\Gamma\leq\mathrm{O}^{+}(L)$ of finite co-volume. Moreover, $\{a_{m},b_{n}\}$ form a set of generators for the ring $M^{!}_{*}(\Gamma)$ of meromorphic modular forms on $\mathscr{D}(L)$ which are holomorphic away from the Heegner divisor $C^{\prime}$.
@@
1.3. Previous work
$S$-polarized K3 surfaces for three cases of $S$ in (1.2) have previously appeared in well-known works of Inose, Matsumoto and other authors. Our work gives another interpretation and proof of their results. Let us briefly explain the geometric motivation behind these examples.
A Kummer surface $\mathcal{Y}=\operatorname{Kum}(E_{1}\times E_{2})$ associated with two non-isogenous elliptic curves $E_{1},E_{2}$ admits several inequivalent elliptic fibrations; see [MR1013073, MR2409557]. In particular, such Kummer surfaces admit an alternate fibration, i.e., an elliptic fibration with section and a Mordell-Weil group that contains a 2-torsion section. Fiber-wise translation by the 2-torsion section gives rise to a canonical symplectic involution, known as van Geemen-Sarti involution. If one factors the Kummer surface $\mathcal{Y}$ by the involution and then resolves the eight occurring singularities, a new K3 surface $\mathcal{X}$ is recovered, related via a rational double-cover map to the Kummer surface. This construction is referred to in the literature as the Nikulin construction. In turn, each surface $\mathcal{X}$ also admits a canonical van Geemen-Sarti involution. Moreover, if one repeats the Nikulin construction on $\mathcal{X}$, the original Kummer surface is recovered, together with a (generically) two-to-one rational map $\pi\colon\mathcal{X}\to\mathcal{Y}$. One can check that the surfaces $\mathcal{X}$ are polarized by the lattice $H\oplus E_{8}(-1)\oplus E_{8}(-1)$; see [MR2369941]. There is also a natural push-forward map $\pi_{*}$ that restricts to a morphism between the transcendental lattices $\mathrm{T}_{\mathcal{X}}$ and $\mathrm{T}_{\mathcal{Y}}$ of $\mathcal{X}$ and $\mathcal{Y}$. In fact, $\pi_{*}$ induces a Hodge isometry between $\mathrm{T}_{\mathcal{X}}(2)$ and $\mathrm{T}_{\mathcal{Y}}$ and we have $\mathrm{T}_{\mathcal{X}}=\mathrm{T}_{E_{1}\times E_{2}}$. This relation is referred to as a Shioda–Inose structure on $\mathcal{X}$.
We shall refer to $\mathcal{X}$ as Inose K3 surfaces as they admit a birational model isomorphic to a projective quartic surface introduced by Inose [MR578868]. Shioda and Inose considered these quartic surfaces as an analogue of Weierstrass equations defining elliptic curves. This analogy is particularly fitting since –as we explained above– the coefficients of the Inose quartic are symmetric expressions of pairs of the modular forms that appear as the coefficients in the Weierstrass equation for an elliptic curve.
The entire picture then generalizes to the Picard rank seventeen case when surfaces $mathcal{Y}$ are principally polarized Kummer surfaces: here, the elliptic fibrations on the Jacobian Kummer surfaces were classified in [MR3263663]. The (generalized) Inose K3 surfaces $\mathcal{X}$ are obtained in a similar manner as before from a unique alternate fibration and polarized by the rank seventeen lattice $H\oplus E_{8}(-1)\oplus E_{7}(-1)$; the details may be found in [MR2427457, MR2824841, MR2935386]. The Inose K3 surfaces $\mathcal{X}$ can again be viewed as K3 surfaces admitting Shioda-Inose structures; see [MR728142, MR2279280, MR3087091]. Now, the coefficients of the defining equations of the Inose K3 surfaces can be expressed in terms of the even generators of the ring of modular forms that were found by Igusa [MR114819, MR141643, MR168805].
Aspects of this construction were generalized for K3 surfaces of lower Picard rank in [MR2824841, MR2254405, MR4015343, CHM19]. Since there are no algebraic Kummer surfaces of Picard rank lower than seventeen, those needed to be replaced by other K3 surfaces; a suitable choice for Picard number sixteen turned out to be the surfaces $\mathcal{Y}$ obtained as double covers of the projective plane branched over the union of six lines. In this way, the rank-seventeen case is recovered by making the six lines tangent to a common conic. The surfaces $\mathcal{Y}$ are polarized by the lattice $H\oplus D_{6}(-1)\oplus D_{4}(-1)^{\oplus 2}$. This family was studied by Matsumoto [MR1103969, MR1204828]. The moduli of the family are well understood and are related to Abelian fourfolds of Weil type [MR3506391, MR1335243]. Again, one can obtain (generalized) Inose K3 surfaces $\mathcal{X}$ of Picard rank sixteen which are polarized by the lattice $H\oplus E_{8}(-1)\oplus D_{6}(-1)\cong H\oplus E_{7}(-1)\oplus E_{7}(-1)$; see [CHM19].
In this article we determine the modular forms that constitute the coefficients in unique equations for all $S$-polarized K3 surfaces with $S$ given in (1.2), generalizing the explicit relation between modular forms and K3 surfaces with automorphism group $(\mathbb{Z}/2\mathbb{Z})^{2}$ down to Picard rank 13.
@@ 4. Certain algebras of modular forms
be the basic weak Jacobi forms, which generate the $\mathbb{C}[E_{4},E_{6}]$-module of even-weight weak Jacobi forms, as defined in Chapter 9 of [EZ85].
@@
generate the $\mathbb{C}[E_{4},E_{6}]$-module of weak Jacobi forms of lattice index $A_{1}^{\oplus n}$ which are invariant under substitutions $z_{i}\mapsto-z_{i}$, and therefore the Weyl group $W(nA_{1})$.
@@
For $1\leq n\leq 4$, the algebra of modular forms $M_{*}(\mathrm{O}^{+}(L_{n}))$ is generated by the Eisenstein series $E_{4}$ and $E_{6}$ and by the theta lifts
@@
The algebra of meromorphic modular forms $M_{*}^{!}(\mathrm{O}^{+}(L_{5}))$ with poles on $r_{2}^{\perp}$ and its orbit under $\mathrm{O}^{+}(L_{5})$ is generated by the Eisenstein series $E_{4}$ and $E_{6}$ and by the meromorphic theta lifts
@@ 5. Computation of van Geemen–Sarti equations
We now determine the coefficients of five of the van Geemen–Sarti equations (1.4) in terms of the generators of $M_{*}(\mathrm{O}^{+}(L_{n}))$ described in the previous section. The argument is an induction on $n=8-k$, beginning with $n=0$ (such that $K_{n}=\{0\}$ and $L=H\oplus H$); the general K3 surface in the associated family has Picard rank 18.
@@
$\mathrm{res}\colon\ M_{*}(\mathrm{O}^{+}(L_{1}))\longrightarrow M_{*}(\mathrm{O}^{+}(L_{0}))$
which through the interpretations $M_{*}(\mathrm{O}^{+}(L_{1}))=M_{2*}(\mathrm{Sp}_{4}(\mathbb{Z}))$ and $M_{*}(\mathrm{O}^{+}(L_{0}))=\mathrm{Sym}^{2}(M_{*}(\mathrm{SL}_{2}(\mathbb{Z}))$ is the map
@@
$\mathrm{res}\colon\ M_{*}(\mathrm{O}^{+}(L_{2}))\longrightarrow M_{*}(\mathrm{O}^{+}(L_{1})).$